# Mean-Shift Distillation for Diffusion Mode Seeking

### v1

@@ 1 Introduction
Soon after image diffusion (Dhariwal & Nichol, 2021) models exploded in popularity, Poole et al. (2022) introduced the idea of using them for image optimization. Intuitively, this can be expressed as the notion that images more likely to be generated by a diffusion model are “better” in the sense of being more faithful to the data distribution the diffusion model was trained on.
Formally, diffusion models provide a mechanism to sample images $x\in\mathcal{I}$ from some learned distribution $p(x)$. We then have a parameter vector $\vartheta\in\mathcal{P}$, along with an image-generating model $g:\mathcal{P}\rightarrow\mathcal{I}$. Given an initialization $\vartheta^{0}$, we seek to optimize a $\vartheta^{k}$ such that $p(g(\vartheta^{k}))>p(g(\vartheta^{0}))$. We expect this to yield an image $g(\vartheta^{k})$ of higher quality, under the metric the diffusion model is trained for.
@@
Score distillation sampling (SDS) (Poole et al., 2022) attempts to address this problem by offering proxies for the density gradient that are easier to estimate. However, their theoretical properties are not rigorously established, and SDS suffers from significant bias as well as variance, yielding inaccurate gradients. Examining the loss landscape of SDS in Figure 1, we indeed see that not only are the maxima of this function not collocated with the modes of $p(x)$, but even in the simplest cases the loss creates “phantom modes” that are well out of distribution.
Our method offers both better alignment with the distribution and lower variance of the gradient estimate.
@@ 2 Related Work
A surprising connection between mean shift and diffusion emerged from the analysis of the optimal denoising model (Karras et al., 2022; Chen et al., 2024). Since the forward (noising) process can be expressed as successive convolutions with a Gaussian kernel, the intermediate distributions are in fact Gaussian-kernel density estimates of the data distributions, with kernel bandwidth proportional to the time parameter. Therefore in the ODE of the reverse (inference) process, the gradient of the denoiser is theoretically equal to the mean-shift vector with appropriate kernel and bandwidth. Mean-shifting on the IVP time domain does not in fact seek modes of the output distribution, but we take advantage of this knowledge to implement mean shift on that domain.
Further related to the analysis of modes in particular, Karras et al. (2024); Bradley & Nakkiran (2024) suggest that applying classifier-free guidance (CFG) (Ho & Salimans, 2021) to diffusion has the effect of sharpening the modes of the output distribution. This guidance does not explicitly seek modes, but we have found that using CFG synergizes well with both SDS and our methodß.
@@
The tendency of SDS to produce over-smoothened results due to high variance is well documented. A plethora of works have been proposed to mitigate this behavior, e.g. to factorize the gradient to reduce the bias (Hertz et al., 2023; Yu et al., 2024; Katzir et al., 2024; Alldieck et al., 2024), or to replace the uniform noise sampling in SDS with noise obtained by running DDIM inversion Liang et al. (2023); Lukoianov et al. (2024). Wang et al. (2023a) propose a control variate for SDS, Wang et al. (2023b); Xu et al. (2024); Yan et al. (2025) improve diversity of generations, and Wang et al. (2024) alleviate the multi-view inconsistency problem.
@@
(a) Ground truth (b) DDIM $\bigstar$ (c) SDS $\otimes$ (d) Our MSD $\otimes$
$\bullet$  data sample     $\otimes$  without guidance     $\bigstar$  with guidance (CFG=4)     Refer to caption   loss landscape
Figure 1: Mode-seeking simulated in a fractal-like 2D distribution with two (orange, gray) classes, adapted from Karras et al. (2024). We compare the behavior of diffusion sampling (DDIM) to optimization-based diffusion distillation, in a class-conditional setting. With class=orange, (a) Ground truth distribution, (b) DDIM sampling, (c) SDS optimization without guidance, and (d) our MSD optimization without guidance.
@@ 3.1 Motivation
Given a pre-trained diffusion model $\epsilon_{\phi}$, the SDS loss penalizes the KL-divergence of a unimodal Gaussian distribution centered around $x$ and the data distribution $p_{\phi}(z_{t};y,t)$ captured by the frozen diffusion model conditioned on text embeddings $y$. With $x=g(\vartheta)$, an image rendered by $\vartheta$ via a differentiable renderer $g$, Poole et al. (2022) derive the gradient of the loss $\mathcal{L}_{\mathrm{SDS}}$ with respect to $\vartheta$:
@@
To illustrate the pitfalls of SDS, we simulate it in 2D using a small denoising diffusion network (Figure 1). This allows us to set $\vartheta=x\in{\mathbb{R}}^{2}$ (where $g$ becomes an identity map). We construct a fractal-like dataset as shown by Karras et al. (2024), with analytic ground-truth probability density and score. This data distribution is a mixture of highly anisotropic Gaussians, where most of the probability mass resides in narrow regions, emulating the low intrinsic dimensionality of natural images (Roweis & Saul, 2000; Belkin & Niyogi, 2003). For a baseline, we compare it with DDIM (Song et al., 2021a), a popular first-order sampling algorithm, with classifier-free guidance (CFG) (Ho & Salimans, 2021). More details can be found in Section 4.2.
@@ 3.3 Gradient Estimation via Product Sampling
$x^{\prime}=\frac{\sum_{y_{i}\sim p}K_{\lambda}(x-y_{i})y_{i}}{\sum_{y_{i}\sim p}K_{\lambda}(x-y_{i})}.$ (7)
@@
which is the sum of the density score (provided by the diffusion model) and the score of our Gaussian kernel.
Having the ability to generate samples $y_{i}$ from the product density, we can estimate the mean-shift iterate (8) as
@@
We can thus step along $\vec{m}$ to seek the modes of the data density $p$. Substituting a learned score model into Equation 9 gives us
@@
$\otimes$  without guidance     $\bigstar$  with guidance     $\blacklozenge$ guidance in limited interval     $f^{-1}$  noise scaling
Figure 2: We juxtapose diffusion sampling vs diffusion distillation in low-dimensional (${\mathbb{R}}^{2}$) and high-dimensional (${\mathbb{R}}^{64\times 64\times 4}$) setting. Top: (a) text-conditioned generation of image via DDIM with 32 steps, (b) - (d) optimized coordinate-based neural implicit image for SDS, SDI, and our MSD respectively with StableDiffusion (CFG=7.5, § 4.4). Bottom: (a) class-conditioned generation of 2D points via DDIM with 32 steps, (b) - (d) optimized 2D points for SDS, SDI, and our MSD respectively (CFG=4, § 4.2). Text-prompts in clockwise order: “A DSLR photo of a … hamburger, squirrel dressed as a samurai weighing a katana, knight in silver armor, and bluejay on basket of macarons".
@@ 3.4 Practical Considerations
We may directly substitute this for the denoiser term in Equation 12. Despite its practical success, the denoising direction induced by CFG does not provide theoretical guarantees in producing samples from $p_{0,w}(z_{t}|c)$ (Bradley & Nakkiran, 2024). Even in simple settings, as observed in Figure 1(b), CFG can lead to mode drops. While alternative guidance strategy exists (Karras et al., 2024), we stick with the dominant practice of using CFG (Equation 13). We have found that this synergizes well with mode-seeking by mean-shift, and show the effects of this in evaluation below.
@@ 4 Practical Implementation and Evaluation
In this section, we construct synthetic examples on which we demonstrate that our proposed method behaves as theory predicts, alleviating the issues SDS exhibits even in these simple scenarios. We further explain the issues encountered when translating this theory into practice, and describe adaptations we designed to make our method work with real-world diffusion models. Finally, we evaluate that in these real-world scenarios, we retain the desirable properties and performance of our method. When applicable, we make comparisons with SDI (Lukoianov et al., 2024), a method that improves upon SDS by proposing a better noise term, yet, retaining the same gradient computation.
@@ 4.1 Idealized Setting
In order to manage large data dimensionality as well as massive training datasets, diffusion in practice employs a trained neural network to represent the denoiser $D$. However, Karras et al. (2022) have identified an analytical solution to minimizing the denoiser error, the ideal denoiser $D^{*}(x;t)$:
$D^{*}(x;t)=\frac{\sum_{i}y_{i}\mathcal{N}(x;y_{i},\sigma(t))}{\sum_{i}\mathcal{N}(x;y_{i},\sigma(t))},$ (16)
where $y_{0}\dots y_{n}$ are samples in our training set. Attentive readers will notice that this is in fact the discrete mean shift formula (Comaniciu & Meer, 2002), with training samples taking the place of data samples and noise magnitude $\sigma(t)$ taking place of the kernel bandwidth $\lambda$. This expression is feasible to compute in practice for small datasets, and we may substitute it into the SDS formula (1) to get an explicit solution for the SDS gradient:
@@
This gives us an integral we can numerically evaluate with reasonable guarantees of its behavior. This allows us to compare both methods on synthetic datasets, eliminating any error introduced by training and evaluating a neural model to show that the theoretical properties hold.
@@ 4.2 Toy Distributions in ${\mathbb{R}}^{2}$
In addition to the fractal dataset (Figure 1), we extend our analysis to other 2D datasets, with ${y_{i}}^{m}_{i=1}\subset{\mathcal{M}}\in{\mathbb{R}}^{2}$ sampled from various challenging toy 2D densities (van der Walt et al., 2014; Rozen et al., 2021). Given $10^{4}$ samples from the data distribution and points densely initialized across a grid $[-1.5,1.5]^{2}$, we evaluate SDS (Algorithm 1) and our MSD (Algorithm 2) with $10^{3}$ Monte Carlo samples. We perform distillation with a learned denoiser (13) and an ideal denoiser (16). For the learned denoiser, we use the architecture and training setup used by Karras et al. (2024) and similarly represent the densities as mixtures of Gaussians. We use the Adam optimizer (Kingma & Ba, 2015) and run the optimization procedure for 150 steps with a learning rate of $0.08$. For our MSD, we set an initial bandwidth of $0.316\sim\sqrt{0.1}$ which is linearly decayed over the course of the optimization. For the ideal denoiser, due to it’s high cost requirements in time and memory, we instead opt to use a few steps of gradient descent with high learning rate.
We visualize the generated samples produced by our method and SDS after the optimization in Figure 3. In addition to optimizing samples, we evaluate both SDS and MSD gradients across the domain and then numerically integrate them to reconstruct the loss functions they represent. This makes the bias in SDS particularly obvious, as the peaks of this reconstructed function may be well out of distribution.
We suspect that this bias persists in SDS in higher dimensional settings and is what causes SDS optimized results to be blurry and exhibit other artifacts (top row of Figure 2).
@@
We measure cost as number of invocations of the score model, since that is the typical bottleneck in diffusion. Normalization by the squared norm of $g$ is included to account for the fact that due to bias and scaling, different estimators may converge to gradients of different magnitude, and the normalized MSE then roughly describes the probability of the estimated gradient pointing the “right” way.
We accumulate both MSE and cost over many independent estimations, and average over many values of $x$.
Table 1: Efficiency $\uparrow$
Dataset SDS Ours
Fractal -6.89 7.65
Spiral -7.57 6.32
Pinwheel -6.99 7.08
The result of these efficiency comparisons between our method and SDS is in Table 1 (in log-scale). Although getting a single estimate with our method requires more score model invocations, the efficiency of our method is significantly higher. We observe similar behavior in complex settings (Figure 4). In Table 2, we evaluate the quality of the generated samples. Our method outperforms SDS, in both the ideal denoiser and learned denoiser setting.
Table 2: Metrics for 2D toy density datasets. Left: ideal denoiser ($D^{*}$) / Right: learned denoiser ($D_{\theta}$). MMD scaled by $10^{-4}$.
Fractal DDIM -1.85 / -1.51 0.97 / 0.95 0.93 / 0.96 0.86 / 0.007
SDS 36.15 / 9.12 0.08 / 0.01 0.03 / 0.0 328.03 / 87.04
Ours -1.32 / -2.02 0.92 / 0.97 0.33 / 0.42 30.46 / 12.79
Spiral DDIM -1.39 / -1.32 0.97 / 0.96 0.93 / 0.96 0.41 / 1.16
SDS 30.37 / 8.13 0.02 / 0.04 0.03 / 0.11 13.85 / 274.35
Ours -1.28 / -1.51 0.99 / 0.98 0.18 / 0.18 4.49 / 18.41
Pinwheel DDIM -1.19 / -1.1 0.97 / 0.97 0.94 / 0.97 1.05 / 0.27
SDS 2.29 / 2.00 0.85 / 0.90 0.03 / 0.005 5.18 / 36.37
Ours -1.94 / -2.19 0.99 / 0.99 0.01 / 0.13 5.83 / 7.25
Figure 3: Unconditional distillation on two toy density datasets, given an ideal denoiser ($D^{*}$) and a learned denoiser ($D_{\theta}$). For each method and denoiser, we show the optimized samples (left) and the loss landscape (right). Zoom in for clarity.

### v2

@@ 1 Introduction
Soon after image diffusion (Dhariwal & Nichol, 2021) models exploded in popularity, DreamFusion (Poole et al., 2022) and SJC (Wang et al., 2022) concurrently introduced the idea of using them for image optimization. Intuitively, this can be expressed as the notion that images more likely to be generated by a diffusion model are “better” in the sense of being more faithful to the data distribution the diffusion model was trained on.
Formally, diffusion models provide a mechanism to sample images $x\in\mathcal{I}$ from a learned distribution $p(x)$. We then have a parameter vector $\vartheta\in\mathcal{P}$, along with an image-generating model $g:\mathcal{P}\rightarrow\mathcal{I}$. Given an initialization $\vartheta^{0}$, we seek to optimize a $\vartheta^{k}$ such that $p(g(\vartheta^{k}))>p(g(\vartheta^{0}))$. We expect this to yield an image $g(\vartheta^{k})$ of higher quality, under the metric the diffusion model is trained for.
@@
Score distillation sampling (SDS) (Poole et al., 2022; Wang et al., 2022) attempts to address this problem by offering proxies for the density gradient that are easier to estimate. However, their theoretical properties are not rigorously established, and SDS suffers from significant bias as well as variance, yielding inaccurate gradients. Examining the loss landscape of SDS in Figure 1, we indeed see that not only are the maxima of this function not collocated with the modes of $p(x)$, but even in the simplest cases the loss creates “phantom modes” that are well out of distribution. Our method offers both better alignment with the distribution and lower variance of the gradient estimate.
@@ 2 Related Work
A surprising connection between mean shift and diffusion emerged from the analysis of the optimal denoising model (Karras et al., 2022; Chen et al., 2024; Je et al., 2024). Since the forward (noising) process can be expressed as successive convolutions with a Gaussian kernel, the intermediate distributions are in fact Gaussian-kernel density estimates of the data distributions, with kernel bandwidth proportional to the time parameter. Therefore in the ODE of the reverse (inference) process, the gradient of the denoiser is theoretically equal to the mean-shift vector with appropriate kernel and bandwidth. Mean-shifting on the IVP time domain does not in fact seek modes of the output distribution, but we take advantage of this knowledge to implement mean shift on that domain.
Further related to the analysis of modes in particular, (Karras et al., 2024; Bradley & Nakkiran, 2024) suggest that applying classifier-free guidance (CFG) (Ho & Salimans, 2021) to diffusion has the effect of sharpening the modes of the output distribution. This guidance does not explicitly seek modes, but we have found that using CFG synergizes well with both SDS and our method.
@@
The tendency of SDS to produce over-smoothened results due to high variance is well documented. A plethora of works have been proposed to mitigate this behavior, e.g. to factorize the gradient to reduce the bias (Hertz et al., 2023; Yu et al., 2024; Katzir et al., 2024; Alldieck et al., 2024), or to replace the uniform noise sampling in SDS with noise obtained by running DDIM inversion (Liang et al., 2023; Lukoianov et al., 2024). SteinDreamer (Wang et al., 2023a) propose a control variate for SDS, (Xu et al., 2024; Yan et al., 2025) improve diversity of generations, and (Wang et al., 2024) alleviate the multi-view inconsistency problem.
VSD (Wang et al., 2023b) tackle the low-fidelity problem by treating the target parameters as a random variable and estimate the variational distribution to produce diverse, high-fidelity results. SDI (Lukoianov et al., 2024), on the other hand and unlike previous variance-reduction methods, find a better approximation of the added noise term, eliminating one of the root causes of the excessive variance. These methods attribute SDS/SJC to be mode-seeking; we show it is not and introduce mode-seeking behavior to address the excessive variance and low-fidelity results.
We draw the distinction between the above methods to knowledge distillation works designed for one-step inference (Yin et al., 2024; Luo et al., 2023; Xie et al., 2024). Diff-Instruct (Luo et al., 2023) show that SDS is a special case of their distillation formulation when the generator’s output is a Dirac’s Delta distribution and the marginal of the variational score is Gaussian. While the derived gradients resemble SDS and VSD, these methods require training an auxiliary score network to estimate the variational score— challenging to generalize beyond text-to-image generation— and have been targeted for different use cases, namely faster inference, model compression, and dataset privacy-preserving.
@@
(a) Ground truth (b) DDIM (c) SDS (d) Our MSD
$\bullet$  data sample     Refer to caption   loss landscape
Figure 1: Mode-seeking simulated in a fractal-like 2D distribution with two (orange, gray) classes, adapted from (Karras et al., 2024). We compare the behavior of diffusion sampling (DDIM) to optimization-based diffusion distillation, in a class-conditional setting. With class=orange, (a) Ground truth distribution, (b) DDIM sampling , (c) SDS, and (d) our MSD. All methods are run without guidance.
@@ 3.1 Motivation
Given a pre-trained diffusion model $\epsilon_{\phi}$, the SDS loss penalizes the KL-divergence of a unimodal Gaussian distribution centered around $x$ and the data distribution $p_{\phi}(z_{t};c,t)$ captured by the frozen diffusion model conditioned on text embeddings $c$. With $x=g(\vartheta)$, an image rendered by $\vartheta$ via a differentiable renderer $g$, (Poole et al., 2022) derive the gradient of the loss $\mathcal{L}_{\mathrm{SDS}}$ with respect to $\vartheta$:
@@
To illustrate the pitfalls of SDS, we simulate it in 2D using a small denoising diffusion network (Figure 1). This allows us to set $\vartheta=x\in{\mathbb{R}}^{2}$ (where $g$ becomes an identity map). We construct a fractal-like dataset as shown by (Karras et al., 2024), with analytic ground-truth probability density and score. This data distribution is a mixture of highly anisotropic Gaussians, where most of the probability mass resides in narrow regions, emulating the low intrinsic dimensionality of natural images (Roweis & Saul, 2000; Belkin & Niyogi, 2003). For a baseline, we compare it with DDIM (Song et al., 2021a), a popular first-order sampling algorithm, with classifier-free guidance (CFG) (Ho & Salimans, 2021). More details can be found in Section 4.2.
@@ 3.3 Gradient Estimation via Product Sampling
$x^{\prime}=\frac{\sum_{y_{i}\sim p}G_{\lambda}(x-y_{i})y_{i}}{\sum_{y_{i}\sim p}G_{\lambda}(x-y_{i})}.$ (7)
@@
where $z_{t}=\alpha(t)x+\epsilon$. This is the sum of the density score (provided by the diffusion model) and the score of our Gaussian kernel. Having the ability to generate samples $y_{i}$ from the product density, we can estimate the mean-shift iterate (8) as
@@
We can thus step along $\vec{m}$ to seek the modes of the data density $p$. Substituting a learned score model into 9 gives us
@@
Figure 2: We juxtapose diffusion sampling vs diffusion distillation in low-dimensional (${\mathbb{R}}^{2}$) and high-dimensional (${\mathbb{R}}^{64\times 64\times 4}$) setting, using guidance via CFG (Ho & Salimans, 2021). Top: (a) text-conditioned generation of image via DDIM with 32 steps, (b) - (e) optimized coordinate-based neural implicit image for SDS, VSD, SDI, and our MSD respectively with StableDiffusion (CFG=7.5, § 4.4). Bottom: (a) class-conditioned generation of 2D points via DDIM with 32 steps, (b) - (e) optimized 2D points for SDS, VSD, SDI, and our MSD respectively (CFG=4, § 4.2). Text-prompts in clockwise order: “A DSLR photo of a … hamburger, squirrel dressed as a samurai weighing a katana, knight in silver armor, and bluejay on basket of macarons".
@@ 3.4 Practical Considerations
As noted in SJC (Wang et al., 2022), distillation with unconditioned diffusion models is challenging in high-dimensional settings like images. While we show unconditioned diffusion distillation is practical in simple 2D toy datasets below, we operate in the conditional setting throughout.
@@
We may directly substitute this for the denoiser term in Equation 12. Despite its practical success, the denoising direction induced by CFG does not provide theoretical guarantees in producing samples from $p_{0,w}(z_{t}|c)$ (Bradley & Nakkiran, 2024). Even in simple settings, as observed in Figure 1(b), CFG can lead to mode drops. While alternative guidance strategy exists (Karras et al., 2024), we stick with the dominant practice of using CFG (Equation 13). We have found that this synergizes well with mode-seeking by mean-shift, and show the effects of this in evaluation below. See discussion in Appendix B.
@@
Table 1: Metrics for class-conditional distillation on 2D Fractal dataset. For each metric, left to right: ideal denoiser ($D^{*}$), learned denoiser ($D_{\theta}$) without guidance, learned denoiser with CFG (Ho & Salimans, 2021), and learned denoiser with Autoguidance (Karras et al., 2024). $1^{\text{st}}$ and $2^{\text{nd}}$ best among distillation-based methods for each column, highlighted.
Method NLL $\downarrow$
DDIM -1.85 -1.51 -1.59 -1.67
SDS 36.15 9.12 15.96 11.33
VSD 9.97 9.88 18.97 11.52
SDI 24.28 -2.87 27.37 0.65
Ours -1.32 -2.02 -1.15 -1.99
Method MMD $\downarrow$ (scaled by $10^{-4}$)
DDIM 0.860 0.007 257.43 0.25
SDS 328.0 87.04 3875.11 71.05
VSD 230.9 94.68 3845.41 70.25
SDI 29.93 459.9 69.23 15089
Ours 30.46 12.79 133.41 122.94
Method Precision $\uparrow$
DDIM 0.97 0.95 0.97 0.96
SDS 0.08 0.01 0.17 0.04
VSD 0.05 0.10 0.21 0.03
SDI 0.27 0.97 0.30 0.51
Ours 0.92 0.97 0.94 0.97
Method Recall $\uparrow$
DDIM 0.93 0.96 0.44 0.79
SDS 0.03 0.00 0.03 0.03
VSD 0.02 0.05 0.05 0.03
SDI 0.01 0.12 0.48 0.51
Ours 0.33 0.42 0.40 0.43
Table 2: Metrics for unconditional distillation on 2D toy datasets. For each metric, left to right: ideal denoiser ($D^{*}$) and learned denoiser ($D_{\theta}$). MMD scaled by $10^{-4}$.
Spiral DDIM -1.39 -1.32 0.97 0.96 0.93 0.96 0.410 1.160
SDS 30.37 8.13 0.02 0.04 0.03 0.11 13.85 274.3
VSD 10.15 8.90 0.04 0.07 0.09 0.14 23.46 271.8
SDI 35.64 19.16 0.10 0.12 0.90 0.42 39.51 2008
Ours -1.28 -1.51 0.99 0.98 0.18 0.18 4.490 18.41
Pinwheel DDIM -1.19 -1.1 0.97 0.97 0.94 0.97 1.05 0.270
SDS 2.29 2.00 0.85 0.90 0.03 0.01 5.18 36.37
VSD 3.34 2.28 0.65 0.97 0.04 0.02 6.78 33.36
SDI 28.31 17.33 0.17 0.51 0.001 0.15 6.13 98.09
Ours -1.94 -2.19 0.99 0.99 0.01 0.13 5.83 7.250
@@ 4 Practical Implementation and Evaluation
In this section, we construct synthetic examples on which we demonstrate that our proposed method behaves as theory predicts, alleviating the issues SDS exhibits even in these simple scenarios. We further explain the issues encountered when translating this theory into practice, and describe adaptations we designed to make our method work with real-world diffusion models, retaining desirable properties. We make comparisons with two strong baselines that improve the convergence and performance of SDS; SDI (Lukoianov et al., 2024), who propose a better noise term to reduce late-stage stochasticity, yet, retain the same gradient computation of SDS, and VSD (Wang et al., 2023b), who propose to learn the variational score as opposed to assuming it to be a known analytic score like in SDS.
@@ 4.1 Idealized Setting
In order to manage large data dimensionality as well as massive training datasets, diffusion in practice employs a trained neural network to represent the denoiser $D$. However, (Karras et al., 2022; Je et al., 2024) have identified an analytical solution to minimizing the denoiser error, the ideal denoiser $D^{*}(x;t)$:
$D^{*}(x;t)=\frac{\sum_{i}u_{i}\mathcal{N}(x;u_{i},\sigma(t))}{\sum_{i}\mathcal{N}(x;u_{i},\sigma(t))},$ (16)
where $u_{0}\dots u_{n}$ are samples in our training set. Attentive readers will notice that this is in fact the discrete mean shift formula (Comaniciu & Meer, 2002), with training samples taking the place of data samples and noise magnitude $\sigma(t)$ taking place of the kernel bandwidth $\lambda$. This expression is feasible to compute in practice for small datasets, and by setting
$\!\!\!\!\epsilon^{*}_{\phi}(z_{t};t)=-\frac{D^{*}(z_{t};t)-z_{t}}{\sigma_{t}},$
we may substitute it into the SDS formula (1) to get an explicit solution for the SDS gradient
@@
We can brute force numerically evaluate this integral. We can compare both methods on synthetic datasets, eliminating any error introduced by training and evaluating a neural model to show that the theoretical properties hold.
Figure 3: Unconditional distillation on two toy density datasets, Pinwheel (top) and Spiral (bottom), given an ideal denoiser ($D^{*}$) and a learned denoiser ($D_{\theta}$). For each method and denoiser, we show the optimized samples (left) and the loss landscape (right). Zoom in for clarity.
@@ 4.2 Toy Distributions in ${\mathbb{R}}^{2}$
In addition to the fractal dataset (Figure 1), we extend our analysis to other 2D datasets, with ${u_{i}}^{m}_{i=1}\subset{\mathcal{M}}\in{\mathbb{R}}^{2}$ sampled from various challenging toy 2D densities (van der Walt et al., 2014; Rozen et al., 2021). For each, we sample $10^{4}$ points from the data distribution and initialize our target points densely across a grid $[-1.5,1.5]^{2}$. With $10^{3}$ Monte Carlo samples, we benchmark SDS, VSD, SDI, and our MSD (Algorithm 2) using both an ideal denoiser (16) and a learned denoiser (13). For the fractal dataset the denoiser is class-conditioned; its score is either left unchanged (without guidance) or guided via CFG or Autoguidance. For the other datasets, the denoiser is unconditioned.
We visualize the generated samples produced by all methods after the optimization in Figures 2 and 3. We also visualize the reconstructed loss functions. This makes the behavior of all methods particularly obvious; the peaks of this reconstructed function are out of distribution for all methods except our MSD. Numerical evaluations in Tables 1 and 2 show our method outperform baselines. We suspect that this bias persists in SDS in higher dimensional settings and is what causes SDS optimized results to be blurry and exhibit other artifacts (top row of Figure 2).
For the learned denoiser, we use the architecture and training setup used by (Karras et al., 2024) and similarly represent the densities as mixtures of Gaussians. We use the Adam optimizer (Kingma & Ba, 2015) and run the optimization procedure for 150 steps with a learning rate of $0.08$. For our MSD, we set an initial bandwidth of $0.316\sim\sqrt{0.1}$ which is linearly decayed over the course of the optimization. For the ideal denoiser, due to it’s high cost requirements in time and memory, we instead opt to use a few steps of gradient descent with high learning rate.
In addition to optimizing samples, we evaluate both SDS, VSD, SDI, and our MSD gradients across the domain and then numerically integrate them to reconstruct the loss functions they represent.
@@
We measure cost as number of invocations of the score model, since that is the typical bottleneck in diffusion.
Normalization by the squared norm of $g$ is included to account for the fact that, due to bias and scaling, different estimators may converge to gradients of different magnitude, and the normalized MSE then roughly describes the probability of the estimated gradient pointing the “right” way. MSE and cost are accumulated over many independent estimations, and average over many values of $x$.
The result of these efficiency comparisons are in Table 3 (in log-scale). Although getting a single estimate with our method requires more score model invocations, the efficiency of our method is significantly higher than SDS and VSD, and comparable to SDI.